### v2

# Azumaya algebras and Barr’s Theorem

Azumaya algebras and Barr’s Theorem
Thierry Coquand Computer Science and Engineering Department, University of Gothenburg, Sweden, coquand@chalmers.se Henri Lombardi Université Marie et Louis Pasteur, CNRS, LmB, FR-25000 Besançon, France, henri.lombardi@univ-fcomte.fr, stefan.neuwirth@univ-fcomte.fr Stefan Neuwirth Université Marie et Louis Pasteur, CNRS, LmB, FR-25000 Besançon, France, henri.lombardi@univ-fcomte.fr, stefan.neuwirth@univ-fcomte.fr
(May 17, 2025)
Introduction
The goal of this paper is to show how one can use the notion of classifying topos of a coherent theory [Joh, CLR, Wraith], here the theory of strictly Henselian local rings [Wraith], to prove some basic results about Azumaya algebras in a constructive setting.
A fundamental invariant associated to any commutative field $F$ is its Brauer group, whose elements are division algebras over $F$, up to isomorphism. Alternatively an element of the Brauer group can be described as a central simple algebra, up to Morita equivalence, and the group operation is then the tensor product. Two important characterisations of the notion of central simple algebra are the following:
  1. 1.
    an algebra $A$ over $F$ which is finite and such that the canonical map $A\otimes_{F}A^{op}\rightarrow End_{F}(A)$ is an isomorphism; or, equivalently,
  2. 2.
    an algebra $A$ which becomes a matrix algebra over an algebraic (resp. separable) extension of $F$.
This notion of central simple algebra over a field has been generalized to the notion of Azumaya algebra over an arbitrary commutative ring $R$. The definition is a generalisation of the first characterisation: an algebra $A$ which is projective of finite type over $R$ and such that the canonical map $A\otimes_{F}A^{op}\rightarrow End_{R}(A)$ is an isomorphism. In this note, we show in a constructive setting that this definition is equivalent to a suitable generalisation of the second characterisation, using a constructively valid version of Barr’s Theorem [Joh].
1 Some coherent theories
We will consider some coherent theories [Joh] in the language of rings.
1.1 Theory of local rings
The theory of local rings can be formulated in a coherent way with the axioms
$0\neq 1~{}~{}~{}~{}~{}~{}~{}~{}~{}\mathsf{inv}(x)\vee\mathsf{inv}(1-x)$
where $\mathsf{inv}(x)$ denotes $\exists_{y}~{}1=xy$.
If we add to our language a new predicate $\mathsf{m}(x)$ with the axioms
$\mathsf{inv}(x)\vee\mathsf{m}(x)~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\neg(\mathsf{m}(x)\wedge\mathsf{inv}(x))$
then we get the theory of local rings that are residually discrete, i.e. where the property of being invertible is now decidable, and the predicate $\mathsf{m}(x)$ corresponds to the unique maximal ideal of the ring. In this theory, we can express in a coherent way that the ring is Henselian by stating that for any monic polynomial $P$, if we have $u$ such that $P(u)\in\mathsf{m}$ and $P^{\prime}(u)\notin\mathsf{m}$, we can find $\alpha$ such that $P(\alpha)=0$ and $\alpha-u\in\mathsf{m}$. Equivalently [ALP], any monic polynomial of degree $l+1$ which is of the form $X^{l}(X-1)$ mod. $\mathsf{m}$ has a root which is equal to $1$ mod. $\mathsf{m}$. The work [ALP] shows in an intuitionistic setting for this theory the lifting property for idempotents of a finitely generated algebra over a Henselian local ring.
It does not seem possible however to express the property of being Henselian in a coherent way in the language of rings (i.e. without introducing the predicate $\mathsf{m}(x)$). Surprisingly, as noticed by G. Wraith [Wraith], the property of being strictly Henselian, corresponding classically to the property of being Henselian with the residue field being separably closed, can be expressed in a coherent way. In this note, we present some remarks connected to this observation. The first one is a small variation of Wraith’s axiomatisation. Both axiomatisations can be used to define the étale site over a ring. Finally, we show in a constructive setting that Azumaya algebras over a ring are exactly the algebras that are locally matrix algebras for the étale topology.
1.2 Theory of separably closed local rings
If $R$ is a ring, we define the Zariski spectrum of $R$ to be the distributive lattice generated by symbols $D(a)$ and relations
$D(1)=1~{}~{}~{}~{}~{}D(0)=0~{}~{}~{}~{}~{}~{}D(a+b)\leqslant D(a)\vee D(b)~{}~{}~{}~{}~{}~{}~{}~{}D(ab)=D(a)\wedge D(b)$
We write $D(a_{1},\dots,a_{n})$ for $D(a_{1})\vee\dots\vee D(a_{n})$.
One way to build this lattice is to define $D(a_{1},\dots,a_{n})$ to be the radical of the ideal $(a_{1},\dots,a_{n})$.
In particular, this shows that $D(a_{1},\dots,a_{n})=1$ if, and only if, $(a_{1},\dots,a_{n})=1$.
It is also direct to see that $D(a,b)=D(a+b,ab)$ and more generally that
$D(a_{1},\dots,a_{n})=D(s_{1},\dots,s_{n})\text{,}$
where $s_{1},\dots,s_{n}$ are the elementary symmetric functions of $a_{1},\dots,a_{n}$.
For the formulation of this theory of strictly Henselian rings, we need the notion of universal decomposition algebra $L=R[x_{1},\dots,x_{n}]=R[X_{1},\dots,X_{n}]/I$ of a monic polynomial
$f=X^{n}-u_{1}X^{n-1}+u_{2}X^{n-2}-\dots$
in $R[X]$, where $I$ is the ideal generated by $u_{1}-\sigma_{1},~{}\dots,~{}u_{n}-\sigma_{n}$, where $\sigma_{1},\dots,\sigma_{n}$ in $R[X_{1},\dots,X_{n}]$ are the elementary symmetric polynomials. We know that $L=R[x_{1},\dots,x_{n}]$ is freely generated as a module over $R$ and it follows from this that $L$ is a faithfully flat extension of $R$.
We define $\delta_{i}(f)$ to be the element in $R$ equal to $\sigma_{i}(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))$ and $\Delta(f)$ to be $D(\delta_{1}(f),\dots,\delta_{n}(f))$.
We say that $f$ is unramifiable if, and only if, $1=\Delta(f)$. Since $D(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))=D(\delta_{1}(f),\dots,\delta_{n}(f))$, this condition is equivalent to $(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))=1$ in $L$.
Over a residually discrete local ring with residue field $F$, to be unramifiable means that the polynomial $f$ has a simple root in some extension of $F$.
Lemma 1.1.
Let $R$ be a local ring residually discrete of maximal ideal $\mathsf{m}$ with residue field $F=R/\mathsf{m}$. If $f$ is a monic polynomial in $R[X]$ which has a simple root in some nontrivial $F$-algebra $A$, then $f$ is unramifiable.
Lemma 1.2.
If $f=(X-a_{1})\dots(X-a_{p})g$ in $R[X]$ with $g$ monic of degree $n$ and $\Delta(f)=1$ and $R$ is local, then $f^{\prime}(a_{1})$ is invertible or $\dots$ or $f^{\prime}(a_{p})$ is invertible or $\Delta(g)=1$.
Proof.
We write $h=(X-a_{1})\cdots(X-a_{p})$.
We have in $R[x_{1},\dots,x_{n}]$, the decomposition algebra of $g$,
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}),h(x_{1})g^{\prime}(x_{1}),\dots,h(x_{n})g^{\prime}(x_{n}))$
and hence
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}),g^{\prime}(x_{1}),\dots,g^{\prime}(x_{n}))$
and from this follows
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}))\vee\Delta(g)$
in $R[x_{1},\dots,x_{n}]$. Since the decomposition algebra is faithfully flat, we get
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}))\vee\Delta(g)$
in $R$, and since $R$ is local we have the conclusion. ∎
Lemma 1.3.
If $R$ is local and any monic unramifiable polynomial has a root in $R$ then any monic unramifiable polynomial has a simple root in $R$.
Proof.
Assume that $R$ is local and any monic unramifiable polynomial has a root in $R$. Let $f$ be a monic unramifiable polynomial. It has a root $a_{1}$ and we can write $f=(X-a_{1})f_{1}$. By Lemma 1.2, $a_{1}$ is a simple root of $f$ or $\Delta(f_{1})=1$. If $\Delta(f_{1})=1$ then it has a root $a_{2}$ and $f=(X-a_{1})(X-a_{2})f_{2}$. Then by Lemma 1.2 again, $a_{1}$ or $a_{2}$ is a simple root of $f$ or $\Delta(f_{2})=1$, and so on until we find a simple root of $f$. ∎
Let $k$ be a ring. We consider now the category of finitely presented $k$-algebras with the covering
  1. 1.
    $S$ is covered by $S\rightarrow S[1/u_{1}],\dots,S\rightarrow S[1/u_{m}]$ if $1=(u_{1},\dots,u_{m})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic and unramifiable.
It also follows from Lemma 1.3 that we can instead take the following covering, since it classifies an equivalent theory.
  1. 1.
    $S$ is covered by $S\rightarrow S[1/u_{1}],\dots,S\rightarrow S[1/u_{m}]$ if $1=(u_{1},\dots,u_{m})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(P)[1/P^{\prime}]$ if $P$ is monic and unramifiable.
This defines a site, in the sense of Grothendieck (see e.g. [Joh]), over the opposite of the category of finitely presented $k$-algebras. The corresponding sheaf model is the classifying topos of the theory of strictly Henselian $k$-algebras [Joh]. In this sheaf model, we have a “generic” $k$-algebra which is strictly Henselian.
We also may state, following Wraith,
Theorem 1.1.
If $R$ is a local ring which is also residually discrete, then $R$ is strictly Henselian if, and only if, $R$ is Henselian and its residue field is separably closed.
Proof.
Let $P$ be a monic polynomial which has residually a simple root $u$. We show by induction on the degree of $P$ that $P$ has a root in $R$ which is equal to $u$ mod. $\mathsf{m}$. The polynomial $P$ is unramifiable by Lemma 1.1, since it has residually a simple root, and it has a root $a$. If $a\neq u$ mod. $\mathsf{m}$ then we write $P=(X-a)Q$, and $u$ is residually a simple root of $Q$. ∎
2 Some lemmas about Henselian rings
In the next section, we need some results about Henselian rings. In order to make the paper self-contained we list here these results with their proofs. They are taken from the reference [ALP], with a variation in the proof of Lemma 2.1.
We assume in this section that $R$ is local residually discrete with maximal ideal $\mathsf{m}$ and Henselian.
Lemma 2.1 ([ALP, Prop. 5.1]).
Let $P$ be a monic polynomial of degree $n$ in $R[X]$ and $R[x]=R[X]/(P)$. If we have $e(X)$ in $R[X]$ such that $e(x)$ is idempotent mod. $\mathsf{m}R[x]$, then we can find $u$ idempotent in $R[x]$ and $u=e(x)$ mod. $\mathsf{m}R[x]$.
Proof.
Let $A=R[x_{1},\dots,x_{n}]$ be the universal decomposition algebra of $P$. This is a free $R$-module over $R$ of dimension $n!$ and the symmetric group acts on it [LQ, Fact 3-4.4]. We may identify $R[x]$ with the subalgebra $R[x_{1}]$ and we have $e(x_{1})$ in $R[x_{1}]$ which is idempotent mod. $\mathsf{m}R[x_{1}]$. We show that we have $u_{1}$ in $R[x_{1}]$ which is idempotent in $R[x_{1}]$ and equal to $e(x_{1})$ mod. $\mathsf{m}R[x_{1}]$.
The product $e(x_{1})\cdots e(x_{n})$ is in $R$ and it is an idempotent mod. $\mathsf{m}$, so it is $1$ or 00 mod. $\mathsf{m}$. If it is $1$, we have $e(x_{1})=1$ mod. $\mathsf{m}R[x_{1}]$ and we can take $u_{1}=1$.
We can thus assume $e(x_{1})\cdots e(x_{n})=0$ mod. $\mathsf{m}$. We write $e_{I}=\prod_{i\in I}e(x_{i})$ for $I$ subset of $1,\dots,n$. We let [^1] The $e_{I},~{}|I|=k$, are pairwise conjugate by the action of the symmetric group and equality is decidable mod. $\mathsf{m}$. $k<n$ be such that $e_{I}\neq 0$ mod. $\mathsf{m}$ if $|I|=k$ and $e_{I}=0$ mod. $\mathsf{m}$ if $|I|=k+1$. Note that the $e_{I}$ for $|I|=k$ are pairwise orthogonal mod. $\mathsf{m}$. We have $\Sigma_{|I|=k}e_{I}=1$ mod. $\mathsf{m}$ since it is an element in $R$, which is idempotent mod. $\mathsf{m}$. Hence the family of the $e_{I},~{}|I|=k$, forms an FSOI [^2] A “Fundamental System of Orthogonal Idempotents” ([LQ, Sec. II-4]). mod. $\mathsf{m}A$.
We deduce that
$Q(X)=\prod_{|I|=k}(X-e_{I})\text{,}$
which is a monic polynomial in $R[X]$, is of the form $X^{N-1}(X-1)$ mod. $\mathsf{m}R[X]$ with $N=\binom{n}{k}$.
Since $R$ is Henselian, we have $\alpha$ in $R$ such that $Q(\alpha)=0$ and $\lambda=Q^{\prime}(\alpha)$ is invertible. If we write $Q_{I}=Q/(X-e_{I})$ then $Q^{\prime}(X)=\Sigma_{I}Q_{I}$. We take
$v_{I}=\frac{1}{\lambda}Q_{I}(\alpha)$
and we then have that the $v_{I},~{}|I|=k$, form an FSOI in $R[x_{1},\dots,x_{n}]$. We have $v_{I}=\Pi_{J\neq I}(1-e_{J})=e_{I}$ mod. $\mathsf{m}A$. Let $L_{1}$ be the set of all $I$ such that $1$ is in $I$ and $|I|=k$. If we write $u_{1}$ for $\Sigma_{I\in L_{1}}v_{I}$, we have $u_{1}$ in $R[x_{1}]$ and $u_{1}$ is equal to $\Sigma_{I\in L_{1}}e_{I}=e(x_{1})$ mod. $\mathsf{m}A$. This equality follows from $e(x_{1})=e(x_{1})(\Sigma_{I}e_{I})=\Sigma_{I}e(x_{1})e_{I}=\Sigma_{I\in L_{1}}e_{I}$ mod. $\mathsf{m}A$. Hence $e(x_{1})-u_{1}\in R[x_{1}]\cap\mathsf{m}A=\mathsf{m}R[x_{1}]$. ∎
For completeness, we recall the statement and proof of Nakayama’s Lemma [LQ, Lemma IX-2.1].
Lemma 2.2.
Let $M$ be a finitely generated $R$-module and $N$ an $R$-submodule such that $M=N+\mathsf{m}M$, then $M=N$.
Proof.
Let $u$ be a vector column of generators of $M$, then we can find a vector $v$ in $N$ and a matrix $P$ with elements in $\mathsf{m}$ such that $u=v+Pu$. We then have $(I-P)u=v$ and hence the elements of $v$ are generators of $M$ since $\det(I-P)$ is invertible. ∎
Corollary 2.1.
Let $A$ be an $R$-algebra which is finitely generated as an $R$-module, and $u$ in $A$ which is invertible mod. $\mathsf{m}A$, then $u$ is invertible in $A$.
Proof.
We have $A=uA+\mathsf{m}A$ and we can apply the previous Lemma. ∎
Lemma 2.3 ([ALP, Prop. 5.8]).
Let $P$ in $R[X]$ be a monic polynomial in $R[X]$ and assume that we have a decomposition $P=fg$ mod. $\mathsf{m}R[X]$ with $f$ and $g$ monic polynomials coprime mod. $\mathsf{m}$. Then we can find a decomposition $P=FG$ in $R[X]$ with $F=f$ and $G=g$ mod. $\mathsf{m}$.
Proof.
We let $R[x]=R[X]/(P)$. We can find an idempotent $e$ in $R[x]$ mod. $\mathsf{m}R[x]$ such that $(f)=(e),~{}(g)=(1-e)$ mod. $\mathsf{m}R[x]$ [ALP, Prop. 3.11]. Using Lemma 2.1, we may assume $e$ idempotent in $R[x]$. Let $n$ (resp. $m$) be the degree of $f$ (resp. $g$). Then $e,ex,\dots,ex^{m-1}$ and $(1-e),(1-e)x,\dots,(1-e)x^{n-1}$ generate $R[x]$ as an $R$-module mod. $\mathsf{m}$. By Lemma 2.2, they generate the $R$-module $R[x]$. In particular, $ex^{m}$ belongs to the $R$-module generated by $e,ex,\dots,ex^{m-1}$ (since $e$ is idempotent in $R[x]$) and we find in this way $G$ monic of degree $m$ such that $eG(x)=0$ in $R[x]$. Similarly we find $F$ monic of degree $n$ such that $(1-e)F(x)=0$ in $R[x]$. We then have $F(x)G(x)=F(x)eG(x)=0$ and hence $P$ divides $FG$. But they have the same degree, hence $P=FG$. ∎
Theorem 2.1 ([ALP, Thm. 5.9]).
If $A$ is an $R$-algebra which is finitely generated as an $R$-module and $a$ in $A$ is idempotent mod. $\mathsf{m}A$, then we can find $e$ in $A$ idempotent in $A$ and such that $e=a$ mod. $\mathsf{m}A$.
Proof.
We consider the subalgebra $R[a]$ of $A$. Since $A$ is finitely generated as an $R$-module we first find $F$ monic in $R[X]$ such that $F(a)=0$. We decompose $F=X^{n}(1-X)^{m}h$ mod. $\mathsf{m}R[X]$, with $h$ monic and $h(0),~{}h(1)$ not in $\mathsf{m}$. If $n=0$ we have $F(0)=h(0)$ mod. $\mathsf{m}$. Since $F(a)=0$, we get that $a$ is invertible mod. $\mathsf{m}A$, and hence invertible in $A$ using Corollary 2.1, and we can take $e=1$. Similarly, if $m=0$, we get that $1-a$ is invertible and we can take $e=0$.
In the remaining case, we can find using Lemma 2.3 a decomposition in monic polynomials $F=PQH$ in $R[X]$ with $P=X^{n},~{}Q=(1-X)^{m},~{}H=h$ mod. $\mathsf{m}$. We also have $U$ and $V$ in $R[X]$ such that $X(1-X)U+HV=Res(X(1-X),H)=1$ [LQ, Fact 3-7.1], [ALP, Lemma 3.10], and so $H(a)V(a)=1$ mod. $\mathsf{m}A$, and so $H(a)$ is invertible mod. $\mathsf{m}A$ and so invertible in $A$ using Corollary 2.1. Since $F(a)=P(a)Q(a)H(a)=0$, we have $P(a)Q(a)=0$ in $A$. We also have $P(a)=a$ and $Q(a)=1-a$ mod. $\mathsf{m}A$ and so $P(a)+Q(a)=1$ mod. $\mathsf{m}A$ and is also invertible in $A$, by Corollary 2.1, with inverse $\mu$. We can then take $e=\mu P(a)$. ∎
3 Azumaya algebras
An Azumaya algebra $A$ over a ring $k$ is an algebra such that
  1. 1.
    $A$ is projective of finite type;
  2. 2.
    the canonical map $A\otimes_{R}A^{op}\rightarrow End_{k}(A)$ is an isomorphism.
Note that we allow the trivial algebra to be an Azumaya algebra.
We have the following characterisation, which generalizes the result that an algebra is central simple over a field if, and only if, it can be split by a separable extension of this field.
Theorem 3.1.
$A$ is an Azumaya algebra over $k$ if, and only if, we can build a finite tree of root $k$ using the following, where all leaves are matrix algebras:
  1. 1.
    $S$ is covered by $S\rightarrow S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic and unramifiable.
If we have such a tree starting from $k$, notice that each extension $S\rightarrow S[X]/(P)$ for $P$ monic is faithfully flat, and $S\rightarrow\prod_{i}S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$ is also faithfully flat. It follows then that $A$ is finite projective [LQ, Thm. VIII-6.7]. The map $A\otimes_{k}A^{op}\rightarrow End_{k}(A)$ is then a bijection since it becomes a bijection by faithfully flat extension.
The converse follows from the following two results.
Lemma 3.1.
If $A$ is an Azumaya algebra over $R$, which is a local and residually discrete and separably closed $k$-algebra, then $A$ is a matrix algebra.
Proof.
We know that $R$ is Henselian and that $R/\mathsf{m}$ is separably closed by Theorem 1.1.
We simplify the argument in Milne’s course notes ([Milne, Chapter IV, Proposition 1.6]), using the constructive development in [ALP] and [CLN]. If $\mathsf{m}$ is the maximal ideal of $R$, we have that $A$ is split over $R/\mathsf{m}$ ([CLN]), and hence we have a matrix algebra decomposition $e_{ij}$ mod. $\mathsf{m}A$. We can then lift the idempotent $e_{11}$ mod. $\mathsf{m}A$ to some idempotent $e$ in $R$ using Theorem 2.1. We have that $Ae$ is projective and hence free since $R$ is a local ring. The map $A\rightarrow End_{R}(Ae)$ is then an isomorphism, since it is a map between two free $R$-modules of the same dimension, and it is residually an isomorphism. ∎
Corollary 3.1.
If $A$ is an Azumaya algebra over $R$, which is a local and separably closed $k$-algebra, then $A$ is a matrix algebra.
Proof.
For a given Azumaya algebra, the conclusion is coherent, and the hypothesis that $A$ is an Azumaya algebra is formulated in a coherent theory. Since it is proved with the extra hypothesis that $R$ is residually discrete, the syntactic version of Barr’s Theorem [^3] We can actually follow the argument in [Joh] and apply Friedman-Dragalin’s translation [Avi] w.r.t. to the proposition expressing that $A$ is a matrix algebra. (constructively valid) shows that it can be proved without this hypothesis. ∎
Theorem 3.1 follows by interpreting this result in the classifying topos of the theory of strictly Henselian $k$-algebra.
The statement of Theorem 3.1 can be seen as a generalisation of the fact that over a field, a central simple algebra becomes a matrix algebra by a faithfully flat extension [CLN] (classically over an algebraic extension).
We can use such a tree to give arguments by “tree induction”. We give some simple examples.
Lemma 3.2.
Any Azumaya algebra $A$ is central.
Proof.
This means that if $a$ in $A$ satisfies $ax=xa$ for all $x$ in $A$ then $a$ is in $R$. We prove this by “tree induction” using Theorem 3.1. The statement holds on the leaves (it holds for a matrix algebra) and descends to the root. ∎
Lemma 3.3.
If $R$ is a local ring and $A$ a matrix algebra of rank $>0$ [LQ, Def. X-2.2] then any automorphism of $A$ is of the form $x\mapsto axa^{-1}$ for some $a$ in $GL_{n}(A)$.
Proof.
Let $A=End_{R}(R^{n})$ with $v_{1},\dots,v_{n}$ the canonical basis of $R^{n}$ and define $e_{ij}v_{k}=\delta_{jk}v_{i}$. Let $\psi$ be an automorphism of $A$. If $p_{ij}=\psi(e_{ij})$, we have that the $p_{i}=p_{ii}$ form an FSOI and $p_{ij}p_{kl}=\delta_{jk}p_{il}$. Since $R$ is local, each $V_{i}=p_{i}(R^{n})$ is free. Furthermore $R^{n}$ is a direct sum of the $V_{i}$’s; the $V_{i}$’s are pairwise isomorphic since $p_{ji}$ defines an isomorphism between $V_{i}$ and $V_{j}$. So they are all free of rank $1$. Let $w_{1}$ be a basis of $V_{1}$ and $w_{j}=p_{j1}w_{1}$. If $a(v_{i})=w_{i}$ we have $\psi(x)=axa^{-1}$. ∎
Theorem 3.2 (Skolem-Noether).
If $A$ is an Azumaya algebra, of rank $>0$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then $M=\{a\in A~{}|~{}\forall_{x}ax=\psi(x)a\}$ is a projective $k$-module of rank $1$.
Proof.
We use Theorem 3.1. The statement holds on the leaves ($M$ is even free of rank $1$ on the leaves by Lemma 3.2) and descends to the root by tree induction using the local-global principle and the descent principle w.r.t. faithfully flat extension for projective modules ([LQ, Thm. VIII-6.7]). ∎
If $k$ is local, the module $M$ is free of rank $1$ and any generator $u$ of $M$ is invertible (since it is invertible on the leaves) and $\psi$ is the inner automorphism defined by $u$, so we recover the usual form of the Skolem-Noether Theorem (any automorphism of a central simple algebra is an inner automorphism).
Claude Quitté suggested the following refinement.
Theorem 3.3.
If $A$ is an Azumaya algebra, of rank $>0$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then
$M=\{a\in A~{}|~{}\forall_{x}ax=\psi(x)a\}$ and $N=\{a\in A~{}|~{}\forall_{x}a\psi(x)=xa\}$
are projective $k$-modules of rank $1$ and $M\otimes_{k}N\simeq k$ and $MN=k1$. Furthermore, $M$ and $N$ are direct factors in $A$.
Proof.
We have on the leaves $MN\subseteq k1$ and this statement descends to the root by tree induction. This can be used to define a map $M\rightarrow Hom(N,k),~{}x\mapsto(y\mapsto xy)$. This map is an isomorphism on the leaves, and this also descends to the root by tree induction. Finally, $M$ and $N$ are direct factors on the leaves, and this fact descends as well. ∎
4 Theory of algebraically closed local rings
To get the theory of algebraically closed local rings, we add to the theory of local rings the axioms
$\exists_{x}~{}P(x)=0$
for any monic nonconstant polynomial $P$.
Wraith conjectured in [Wraith] that this should define the classifying theory of the fppf topos. This question is discussed further in Blechschmidt’s PhD thesis [BT].
The same argument as in Lemma 3.1 shows the following result.
Lemma 4.1.
If $A$ is an Azumaya algebra over $R$, which is a local and residually discrete and algebraically closed $k$-algebra, then $A$ is a matrix algebra.
We can then prove as in the previous section
Theorem 4.1.
$A$ is an Azumaya algebra over $k$ if, and only if, we can build a finite tree of root $k$ using the following, where all leaves are matrix algebras:
  1. 1.
    $S$ is covered by $S\rightarrow S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic nonconstant.
5 Conclusion
This paper is a beginning of the theory of Azumaya algebras in a constructive setting. The natural next step will be to analyse Gabber’s Theorem [KO] that the group of Azumaya algebras is the torsion part of the second cohomology group with coefficients in $GL_{n}(R)$ for the étale topology.