### v3

# Azumaya algebras and Barr’s Theorem

Azumaya algebras and Barr’s Theorem
Thierry Coquand Henri Lombardi Stefan Neuwirth
(November 24, 2025)
Introduction
The goal of this paper is to show how one can use the notion of classifying topos of a coherent theory [Joh, CLR, Wraith], here the theory of strictly Henselian local rings [Wraith], to prove some basic results about Azumaya algebras in a constructive setting.
A fundamental invariant associated to any commutative field $F$ is its Brauer group, whose elements are division algebras over $F$, up to isomorphism. Alternatively an element of the Brauer group can be described as a central simple algebra, up to Morita equivalence, and the group operation is then the tensor product. Two important characterisations of the notion of central simple algebra are the following [^1] We study these characterisations in a constructive setting in the reference [CLN].:
  1. 1.
    an algebra $A$ over $F$ which is finite and such that the canonical map $A\otimes_{F}A^{op}\rightarrow End_{F}(A)$ is an isomorphism; or, equivalently,
  2. 2.
    an algebra $A$ which becomes a matrix algebra over an algebraic extension of $F$, or, equivalently, which becomes a matrix algebra over an algebraic and separable extension of $F$.
This notion of central simple algebra over a field has been generalized to the notion of Azumaya algebra over an arbitrary commutative ring $R$. The definition is a generalisation of the first characterisation: an algebra $A$ which is projective of finite type over $R$ and such that the canonical map $A\otimes_{F}A^{op}\rightarrow End_{R}(A)$ is an isomorphism. In this note, we show in a constructive setting that this definition is equivalent to a suitable generalisation of the second characterisation, using a constructively valid version of Barr’s Theorem [Joh].
1 Some coherent theories
We will consider some coherent theories [Joh] in the language of rings.
1.1 Theory of local rings
The theory of local rings can be formulated in a coherent way with the axioms
$0\neq 1~~~~~~~~~\mathsf{inv}(x)\vee\mathsf{inv}(1-x)$
where $\mathsf{inv}(x)$ denotes $\exists_{y}~1=xy$.
If we add to our language a new predicate $\mathsf{m}(x)$ with the axioms
$\mathsf{inv}(x)\vee\mathsf{m}(x)~~~~~~~~~~\neg(\mathsf{m}(x)\wedge\mathsf{inv}(x))$
then we get the theory of local rings that are residually discrete [^2] The terminology “discrete” for a set with a decidable equality, while conflicting with some topological intuitions, is well established in constructive mathematics since the work of Bishop [Bish], and we follow this usage in this paper., i.e. where the property of being invertible is now decidable, and the predicate $\mathsf{m}(x)$ corresponds to the unique maximal ideal of the ring. In this theory, we can express in a coherent way that the ring is Henselian by stating that for any monic polynomial $P$, if we have $u$ such that $P(u)\in\mathsf{m}$ and $P^{\prime}(u)\notin\mathsf{m}$, we can find $\alpha$ such that $P(\alpha)=0$ and $\alpha-u\in\mathsf{m}$. Equivalently [ALP], any monic polynomial of degree $l+1$ which is of the form $X^{l}(X-1)$ mod. $\mathsf{m}$ has a root which is equal to $1$ mod. $\mathsf{m}$. The work [ALP] shows in an intuitionistic setting for this theory the lifting property for idempotents of a finitely generated algebra over a Henselian local ring.
It does not seem possible however to express the property of being Henselian in a coherent way in the language of rings (i.e. without introducing the predicate $\mathsf{m}(x)$). Surprisingly, as noticed by G. Wraith [Wraith], the property of being strictly Henselian, corresponding classically to the property of being Henselian with the residue field being separably closed, can be expressed in a coherent way. In this note, we present some remarks connected to this observation. The first one is a small variation of Wraith’s axiomatisation. Both axiomatisations can be used to define the étale site over a ring. Finally, we show in a constructive setting that Azumaya algebras over a ring are exactly the algebras that are locally matrix algebras for the étale topology.
1.2 Theory of separably closed local rings
If $R$ is a ring, and $a_{1},\dots,a_{n}$ elements of $R$, we write $(a_{1},\dots,a_{n})$ for the ideal of $R$ generated by $a_{1},\dots,a_{n}$.
We define, following Joyal [Joy], the Zariski spectrum of $R$ as the universal support $D:R\rightarrow Z(R)$ over $R$. A support is a map $d:R\rightarrow M$ into a bounded distributive lattice $M$ satisfying the relations
$d(1)=1~~~~~d(0)=0~~~~~~d(a+b)\leqslant d(a)\vee d(b)~~~~~~~~d(ab)=d(a)\wedge d(b)$
We write $d(a_{1},\dots,a_{n})$ for $d(a_{1})\vee\dots\vee d(a_{n})$. It is direct that $d(a)\leqslant d(a_{1},\dots,a_{n})$ as soon as $a$ belongs to the radical of the ideal $(a_{1},\dots,a_{n})$. Conversely, it is elementary to check that we get a support if we define $D(a_{1},\dots,a_{n})$ to be the radical of the ideal $(a_{1},\dots,a_{n})$, and this is a concrete way to build the universal support.
In particular, this shows that $D(a_{1},\dots,a_{n})=1$ if, and only if, $(a_{1},\dots,a_{n})=1$.
It is also direct, for any support $d$, to see that $d(a,b)=d(a+b,ab)$ and more generally that
$d(a_{1},\dots,a_{n})=d(s_{1},\dots,s_{n})\text{,}$
where $s_{1},\dots,s_{n}$ are the elementary symmetric functions of $a_{1},\dots,a_{n}$.
For the formulation of this theory of strictly Henselian rings, we need the notion of universal splitting algebra [LQ, III, 4] $L=R[x_{1},\dots,x_{n}]=R[X_{1},\dots,X_{n}]/I$ of a monic polynomial
$f=X^{n}-u_{1}X^{n-1}+u_{2}X^{n-2}$-$\dots$
in $R[X]$, where $I$ is the ideal generated by $u_{1}-\sigma_{1},~\dots,~u_{n}-\sigma_{n}$, where $\sigma_{1},\dots,\sigma_{n}$ in $R[X_{1},\dots,X_{n}]$ are the elementary symmetric polynomials. We know, using Cauchy modules [LQ, III, 4.3], that $L=R[x_{1},\dots,x_{n}]$ is freely generated of rank $n!$ as a module over $R$ and it follows from this that $L$ is a faithfully flat extension of $R$.
For the notion of faithfully flat extension, we follow the definition in [LQ, VIII, 6]: it is an $R$-algebra $i:R\rightarrow S$ such that
  • if $u$ is a row vector in $A^{n}$ and $v$ a column vector in $B^{n}$ such that $i(u)v=0$ then we can find a matrix $M$ with coefficients in $A$ and $w$ such that $uA=0$ and $v=i(A)w$;
  • if $i(a)$ belongs to the ideal $(i(a_{1}),\dots,i(a_{n}))$ in $B$ then $a$ belongs to the ideal $(a_{1},\dots,a_{n})$ in $A$.
It then follows [LQ, VIII, 6] that $i$ is injective. It is also clear from this definition that $S$ is faithfully flat over $R$ when it is a finite free $R$-module of rank >0>0.
We define $\delta_{i}(f)$ to be the element in $R$ equal to $\sigma_{i}(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))$ and $\Delta(f)$ to be $D(\delta_{1}(f),\dots,\delta_{n}(f))$.
We say that $f$ is unramifiable if, and only if, $1=\Delta(f)$. Since $D(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))=D(\delta_{1}(f),\dots,\delta_{n}(f))$, this condition is equivalent to $(f^{\prime}(x_{1}),\dots,f^{\prime}(x_{n}))=1$ in $L$.
Over a residually discrete local ring with residue field $F$, to be unramifiable means that the polynomial $f$ has a simple root in some extension of $F$. This holds in classical mathematics, but also constructively with a dynamical interpretation of extensions of $F$.
Lemma 1.1.
Let $R$ be a local ring, residually discrete with maximal ideal $\mathsf{m}$ and residue field $F=R/\mathsf{m}$. If $f$ is a monic polynomial in $R[X]$ which has a simple root in some nontrivial $F$-algebra AA, then $f$ is unramifiable.
Lemma 1.2.
If $f=(X-a_{1})\dots(X-a_{p})g$ in $R[X]$ with $g$ monic of degree $n$ and $\Delta(f)=1$ and $R$ is local, then $f^{\prime}(a_{1})$ is invertible or $\dots$ or $f^{\prime}(a_{p})$ is invertible or $\Delta(g)=1$.
Proof.
We write $h=(X-a_{1})\cdots(X-a_{p})$.
We have in $R[x_{1},\dots,x_{n}]$, the universal splitting algebra of $g$,
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}),h(x_{1})g^{\prime}(x_{1}),\dots,h(x_{n})g^{\prime}(x_{n}))$
and hence
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}),g^{\prime}(x_{1}),\dots,g^{\prime}(x_{n}))$
and from this follows
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}))\vee\Delta(g)$
in $R[x_{1},\dots,x_{n}]$. Since the universal splitting algebra is faithfully flat, we get
$1=D(f^{\prime}(a_{1}),\dots,f^{\prime}(a_{p}))\vee\Delta(g)$
in $R$, and since $R$ is local we have the conclusion. ∎
Lemma 1.3.
If $R$ is local and any monic unramifiable polynomial has a root in $R$ then any monic unramifiable polynomial has a simple root in $R$.
Proof.
Assume that $R$ is local and any monic unramifiable polynomial has a root in $R$. Let $f$ be a monic unramifiable polynomial. It has a root $a_{1}$ and we can write $f=(X-a_{1})f_{1}$. By Lemma 1.2, $a_{1}$ is a simple root of $f$ or $\Delta(f_{1})=1$. If $\Delta(f_{1})=1$ then it has a root $a_{2}$ and $f=(X-a_{1})(X-a_{2})f_{2}$. Then by Lemma 1.2 again, $a_{1}$ or $a_{2}$ is a simple root of $f$ or $\Delta(f_{2})=1$, and so on until we find a simple root of $f$. ∎
Let $k$ be a ring. We consider now the category of finitely presented $k$-algebras with the covering
  1. 1.
    $S$ is covered by $S\rightarrow S[1/u_{1}],\dots,S\rightarrow S[1/u_{m}]$ if $1=(u_{1},\dots,u_{m})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic and unramifiable.
It also follows from Lemma 1.3 that we can instead take the following covering, since it classifies an equivalent theory.
  1. 1.
    $S$ is covered by $S\rightarrow S[1/u_{1}],\dots,S\rightarrow S[1/u_{m}]$ if $1=(u_{1},\dots,u_{m})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(P)[1/P^{\prime}]$ if $P$ is monic and unramifiable.
This defines a site, in the sense of Grothendieck (see e.g. [Joh]), over the opposite of the category of finitely presented $k$-algebras. The corresponding sheaf model is the classifying topos of the theory of strictly Henselian $k$-algebras [Joh]. In this sheaf model, we have a “generic” $k$-algebra which is strictly Henselian.
We also may state, following Wraith,
Theorem 1.1.
If $R$ is a local ring which is also residually discrete, then $R$ is strictly Henselian if, and only if, $R$ is Henselian and its residue field is separably closed.
Proof.
Let $P$ be a monic polynomial which has residually a simple root $u$. We show by induction on the degree of $P$ that $P$ has a root in $R$ which is equal to $u$ mod. $\mathsf{m}$. The polynomial $P$ is unramifiable by Lemma 1.1, since it has residually a simple root, and it has a root $a$. If $a\neq u$ mod. $\mathsf{m}$ then we write P=(Xa)QP=(X-a)Q, and $u$ is residually a simple root of $Q$. ∎
2 Some lemmas about Henselian rings
In the next section, we need some results about Henselian rings. In order to make the paper self-contained we list here these results with their proofs. They are taken from the reference [ALP], with a variation in the proof of Lemma 2.1.
We assume in this section that $R$ is local residually discrete with maximal ideal $\mathsf{m}$ and Henselian.
Lemma 2.1 ([ALP, Prop. 5.1]).
Let $P$ be a monic polynomial of degree $n$ in $R[X]$ and $R[x]=R[X]/(P)$. If we have $e(X)$ in $R[X]$ such that $e(x)$ is idempotent mod. $\mathsf{m}R[x]$, then we can find $u$ idempotent in $R[x]$ and $u=e(x)$ mod. $\mathsf{m}R[x]$.
Proof.
Let $A=R[x_{1},\dots,x_{n}]$ be the universal splitting algebra of $P$. This is a free $R$-module over $R$ of dimension $n!$ and the symmetric group acts on it [LQ, Fact III-4.4]. We may identify $R[x]$ with the subalgebra $R[x_{1}]$ and we have $e(x_{1})$ in $R[x_{1}]$ which is idempotent mod. $\mathsf{m}R[x_{1}]$. We show that we have $u_{1}$ in $R[x_{1}]$ which is idempotent in $R[x_{1}]$ and equal to $e(x_{1})$ mod. $\mathsf{m}R[x_{1}]$.
The product $e(x_{1})\cdots e(x_{n})$ is in $R$ and it is an idempotent mod. $\mathsf{m}$, so it is $1$ or 0 mod. $\mathsf{m}$. If it is $1$, we have $e(x_{1})=1$ mod. $\mathsf{m}R[x_{1}]$ and we can take $u_{1}=1$.
We can thus assume $e(x_{1})\cdots e(x_{n})=0$ mod. $\mathsf{m}$. We write $e_{I}=\prod_{i\in I}e(x_{i})$ for $I$ subset of $1,\dots,n$. We let [^3] The $e_{I},~|I|=k$, are pairwise conjugate by the action of the symmetric group and equality is decidable mod. $\mathsf{m}$. $k<n$ be such that $e_{I}\neq 0$ mod. $\mathsf{m}$ if $|I|=k$ and $e_{I}=0$ mod. $\mathsf{m}$ if $|I|=k+1$. Note that the $e_{I}$ for $|I|=k$ are pairwise orthogonal mod. $\mathsf{m}$. We have $\Sigma_{|I|=k}e_{I}=1$ mod. $\mathsf{m}$ since it is an element in $R$, which is idempotent mod. $\mathsf{m}$. Hence the family of the $e_{I},~|I|=k$, forms an FSOI [^4] A “Fundamental System of Orthogonal Idempotents” ([LQ, Sec. II-4]). mod. $\mathsf{m}A$.
We deduce that
$Q(X)=\prod_{|I|=k}(X-e_{I})\text{,}$
which is a monic polynomial in $R[X]$, is of the form $X^{N-1}(X-1)$ mod. $\mathsf{m}R[X]$ with $N=\binom{n}{k}$.
Since $R$ is Henselian, we have $\alpha$ in $R$ such that $Q(\alpha)=0$ and $\lambda=Q^{\prime}(\alpha)$ is invertible. If we write $Q_{I}=Q/(X-e_{I})$ then $Q^{\prime}(X)=\Sigma_{I}Q_{I}$. We take
$v_{I}=\frac{1}{\lambda}Q_{I}(\alpha)$
and we then have that the $v_{I},~|I|=k$, form an FSOI in $R[x_{1},\dots,x_{n}]$. We have $v_{I}=\Pi_{J\neq I}(1-e_{J})=e_{I}$ mod. $\mathsf{m}A$. Let $L_{1}$ be the set of all $I$ such that $1$ is in $I$ and $|I|=k$. If we write $u_{1}$ for $\Sigma_{I\in L_{1}}v_{I}$, we have $u_{1}$ in $R[x_{1}]$ and $u_{1}$ is equal to $\Sigma_{I\in L_{1}}e_{I}=e(x_{1})$ mod. $\mathsf{m}A$. This equality follows from $e(x_{1})=e(x_{1})(\Sigma_{I}e_{I})=\Sigma_{I}e(x_{1})e_{I}=\Sigma_{I\in L_{1}}e_{I}$ mod. $\mathsf{m}A$. Hence $e(x_{1})-u_{1}\in R[x_{1}]\cap\mathsf{m}A=\mathsf{m}R[x_{1}]$. ∎
For completeness, we recall the statement and proof of Nakayama’s Lemma [LQ, Lemma IX-2.1].
Lemma 2.2.
Let $M$ be a finitely generated $R$-module and $N$ an $R$-submodule such that $M=N+\mathsf{m}M$, then $M=N$.
Proof.
Let $u$ be a vector column of generators of $M$: then we can find a vector $v$ in $N$ and a matrix $P$ with elements in $\mathsf{m}$ such that $u=v+Pu$. We then have (IP)u=v(I-P)u=v and hence the elements of $v$ are generators of $M$ since $\det(I-P)$ is invertible. ∎
Corollary 2.1.
Let $A$ be an $R$-algebra which is finitely generated as an $R$-module, and $u$ in $A$ which is invertible mod. $\mathsf{m}A$. Then $u$ is invertible in $A$.
Proof.
We have $A=uA+\mathsf{m}A$ and we can apply the previous Lemma. ∎
Lemma 2.3 ([ALP, Prop. 5.8]).
Let $P$ in $R[X]$ be a monic polynomial in $R[X]$ and assume that we have a decomposition $P=fg$ mod. $\mathsf{m}R[X]$ with $f$ and $g$ monic polynomials coprime mod. $\mathsf{m}$. Then we can find a decomposition $P=FG$ in $R[X]$ with $F=f$ and $G=g$ mod. $\mathsf{m}$.
Proof.
We let $R[x]=R[X]/(P)$. We can find an idempotent $e$ in $R[x]$ mod. $\mathsf{m}R[x]$ such that (f)=(e),(g)=(1e)(f)=(e),~(g)=(1-e) mod. $\mathsf{m}R[x]$ [ALP, Prop. 3.11]. Using Lemma 2.1, we may assume $e$ idempotent in $R[x]$. Let $n$ (resp. $m$) be the degree of $f$ (resp. $g$). Then $e,ex,\dots,ex^{m-1}$ and $(1-e),(1-e)x,\dots,(1-e)x^{n-1}$ generate $R[x]$ as an $R$-module mod. $\mathsf{m}$. By Lemma 2.2, they generate the $R$-module $R[x]$. In particular, $ex^{m}$ belongs to the $R$-module generated by $e,ex,\dots,ex^{m-1}$ (since $e$ is idempotent in $R[x]$) and we find in this way $G$ monic of degree $m$ such that $eG(x)=0$ in $R[x]$. Similarly we find $F$ monic of degree $n$ such that (1e)F(x)=0(1-e)F(x)=0 in $R[x]$. We then have $F(x)G(x)=F(x)eG(x)=0$ and hence $P$ divides $FG$. But they have the same degree, hence $P=FG$. ∎
Theorem 2.1 ([ALP, Thm. 5.9]).
If $A$ is an $R$-algebra which is finitely generated as an $R$-module and $a$ in $A$ is idempotent mod. $\mathsf{m}A$, then we can find $e$ in $A$ idempotent in $A$ and such that $e=a$ mod. $\mathsf{m}A$.
Proof.
We consider the subalgebra $R[a]$ of $A$. Since $A$ is finitely generated as an $R$-module we first find $F$ monic in $R[X]$ such that $F(a)=0$. We decompose $F=X^{n}(1-X)^{m}h$ mod. $\mathsf{m}R[X]$, with $h$ monic and $h(0)$, $h(1)$ not in $\mathsf{m}$. If $n=0$ we have $F(0)=h(0)$ mod. $\mathsf{m}$. Since $F(a)=0$, we get that $a$ is invertible mod. $\mathsf{m}A$, and hence invertible in $A$ using Corollary 2.1, and we can take $e=1$. Similarly, if $m=0$, we get that 1a1-a is invertible and we can take $e=0$.
In the remaining case, we can find using Lemma 2.3 a decomposition in monic polynomials $F=PQH$ in $R[X]$ with $P=X^{n},~Q=(1-X)^{m},~H=h$ mod. $\mathsf{m}$. We also have $U$ and $V$ in $R[X]$ such that X(1X)U+HV=Res(X(1X),H)=1X(1-X)U+HV=Res(X(1-X),H)=1 [LQ, Fact 3-7.1], [ALP, Lemma 3.10], and so $H(a)V(a)=1$ mod. $\mathsf{m}A$, and so $H(a)$ is invertible mod. $\mathsf{m}A$ and so invertible in $A$ using Corollary 2.1. Since $F(a)=P(a)Q(a)H(a)=0$, we have $P(a)Q(a)=0$ in $A$. We also have $P(a)=a$ and Q(a)=1aQ(a)=1-a mod. $\mathsf{m}A$ and so $P(a)+Q(a)=1$ mod. $\mathsf{m}A$ and is also invertible in $A$, by Corollary 2.1, with inverse $\mu$. We can then take $e=\mu P(a)$. ∎
3 Azumaya algebras
An Azumaya algebra $A$ over a ring $k$ is an algebra such that
  1. 1.
    AA is projective of finite type;
  2. 2.
    the canonical map $A\otimes_{R}A^{op}\rightarrow End_{k}(A)$ is an isomorphism.
Note that we allow the trivial algebra to be an Azumaya algebra. Note also that any matrix algebra is Azumaya.
One can equivalently describe the algebra structure by a bilinear map $\mu:A\times A\rightarrow A$; its scalar extension along $k\rightarrow S$ is the unique $S$-algebra $A_{S}$ through which every bilinear map extending $\mu$ factors. Thus all subsequent arguments can be phrased without explicit tensor notation.
Over a finite free $k$-module, to give an algebra structure is represented concretely by a multiplication table. A module $A$, projective of finite type over $k$, is given concretely by a projection matrix over $k$ [LQ], and we can similarly represent the algebra structure by a multiplication table (so $m^{3}$ elements of $k$ if the projection matrix has finite size $m$). We can then compute [LQ, X, Theorem 1.6] comaximal elements $s_{1},\dots,s_{l}$ such that $A_{s_{i}}$ is free over $k_{s_{i}}$, and the second condition for being an Azumaya algebra becomes then that some determinants $\delta_{1},\dots,\delta_{l}$ are invertible [^5] After localisation to $s_{j}$ the Azumaya algebra becomes free of rank $r_{i}=n_{i}^{2}\leqslant m$.. From this discussion follows that if $A$ is an Azumaya algebra, given by a multiplication table, and we have a $k$-algebra $k\rightarrow S$ then the algebra $A_{S}$ over $S$, with the same multiplication in $S$, is also an Azumaya algebra.
We have the following characterisation, which generalizes the result that an algebra is central simple over a field if, and only if, it can be split by a separable extension of this field.
Theorem 3.1.
AA is an Azumaya algebra over $k$ if, and only if, we can build a finite tree with root $k$ using the following, where at all leaves all scalar extensions $A_{S}$ are matrix algebras:
  1. 1.
    $S$ is covered by $S\rightarrow S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic and unramifiable.
We prove first that if we have such a tree then the algebra $A$ is Azumaya. If we have such a tree starting from $k$, notice that each extension $S\rightarrow S[X]/(P)$ for $P$ monic is faithfully flat, and $S\rightarrow\prod_{i}S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$ is also faithfully flat. It follows then that $A$ is finite projective [LQ, Thm. VIII-6.7]. The map $A\otimes_{k}A^{op}\rightarrow End_{k}(A)$ is then a bijection since it becomes a bijection by a faithfully flat extension. (As we saw before, to be a bijection is witnessed by the fact that some determinants are invertible, and an element is invertible if it becomes invertible by a faithfully flat extension.)
For the converse, we look at the algebra $A$ (given concretely by its multiplication table) inside the étale topos over $k$, using the site defined in the first section. (This is similar to the constructive use of the algebraic closure in [MC].) The generic ring $R$, defined by $R(S)=S$ for $S$ a finitely presented $k$-algebra, is, in the internal logic of this topos, a local strictly Henselian ring. The algebra $A$ defines then an $R$-algebra $A_{R}(S)=A_{S}$, which is an Azumaya algebra, given by the same multiplication table as the one of $A$. If we can prove in intuitionistic logic that an Azumaya algebra over a strictly Henselian ring is a matrix algebra, we will get, using the interpretation of intuitionistic logic in a sheaf model [Joh, MC], a covering tree of $k$ which gives Theorem 3.1.
Internally in the classifying topos of stricly Henselian $k$-algebras, the object $R$ is local strictly Henselian and the claim that $A_{R}$ is a matrix algebra is a coherent statement, provable intuitionistically. Such a proof yields externally a finite covering tree of $k$ satisfying the conditions above.
The converse is a consequence of the following two results.
Lemma 3.1.
If $A$ is an Azumaya algebra over $R$, which is a local and residually discrete and separably closed $k$-algebra, then $A$ is a matrix algebra.
Proof.
We know that $R$ is Henselian and that $R/\mathsf{m}$ is separably closed by Theorem 1.1.
We simplify the argument in Milne’s course notes ([Milne, Chapter IV, Proposition 1.6]), using the constructive development in [ALP] and [CLN]. If $\mathsf{m}$ is the maximal ideal of $R$, we have that $A$ is split over $R/\mathsf{m}$ ([CLN]), and hence we have a matrix algebra decomposition $e_{ij}$ mod. $\mathsf{m}A$. We can then lift the idempotent $e_{11}$ mod. $\mathsf{m}A$ to some idempotent $e$ in $R$ using Theorem 2.1. We have that $Ae$ is projective and hence free since $R$ is a local ring. The map $A\rightarrow End_{R}(Ae)$ is then an isomorphism, since it is a map between two free $R$-modules of same dimension, and it is residually an isomorphism. ∎
The next proof uses Barr’s Theorem, similar to the way it is used in [Penon, Wraith], but the difference is that we only need a constructively valid version of this theorem.
Corollary 3.1.
If $A$ is an Azumaya algebra over $R$, which is a local and separably closed $k$-algebra, then $A$ is a matrix algebra.
Proof.
For a given Azumaya algebra, the conclusion is coherent [Joh, MC], and the hypothesis that $A$ is an Azumaya algebra is formulated in a coherent theory. Since it is proved with the extra hypothesis that $R$ is residually discrete, the syntactic version of Barr’s Theorem [^6] We can actually follow the argument in [Joh] and apply Friedman-Dragalin’s translation [Avi] with respect to the proposition expressing that $A$ is a matrix algebra. (constructively valid) shows that it can be proved without this hypothesis. ∎
Theorem 3.1 follows by interpreting this result in the classifying topos of the theory of strictly Henselian $k$-algebras.
The statement of Theorem 3.1 can be seen as a generalisation of the fact that over a field, a central simple algebra becomes a matrix algebra by a faithfully flat extension [CLN] (classically over an algebraic extension).
We can use such a tree to give arguments by “tree induction”. We give some simple examples.
Lemma 3.2.
Any Azumaya algebra $A$ is central.
Proof.
This means that if $a$ in $A$ satisfies $ax=xa$ for all $x$ in $A$ then $a$ is in $R$. We prove this by “tree induction” using Theorem 3.1. The statement holds on the leaves (it holds for a matrix algebra) and descends to the root. ∎
Lemma 3.3.
If $R$ is a local ring and $A$ a matrix algebra of rank $\geqslant 1$ [LQ, Def. X-2.2] then any automorphism of $A$ is of the form $x\mapsto axa^{-1}$ for some $a$ in $GL_{n}(A)$.
Proof.
Let $A=End_{R}(R^{n})$ with $v_{1},\dots,v_{n}$ the canonical basis of $R^{n}$ and define $e_{ij}v_{k}=\delta_{jk}v_{i}$. Let $\psi$ be an automorphism of $A$. If $p_{ij}=\psi(e_{ij})$, we have that the $p_{i}=p_{ii}$ form an FSOI and $p_{ij}p_{kl}=\delta_{jk}p_{il}$. Since $R$ is local, each $V_{i}=p_{i}(R^{n})$ is free. Furthermore $R^{n}$ is a direct sum of the $V_{i}$’s; the $V_{i}$’s are pairwise isomorphic since $p_{ji}$ defines an isomorphism between $V_{i}$ and $V_{j}$. So they are all free of rank $1$. Let $w_{1}$ be a basis of $V_{1}$ and $w_{j}=p_{j1}w_{1}$. If $a(v_{i})=w_{i}$ we have $\psi(x)=axa^{-1}$. ∎
For the next result, we use the local-global principle and the descent principle with respect to faithfully flat extension for projective modules ([LQ, Theorem VIII-6.7]). For completeness, we recall the statement. Note that the proof in [LQ] is constructive, using the concrete definition of faithfully flat extension and the fact that, also constructively, a module is finite projective if, and only if, it is finitely presented and flat.
Theorem 3.2 (Descent).
Let $A\rightarrow B$ be a faithfully flat extension and $M$ an $A$-module. If $M\otimes_{A}B$ is finite projective, then so is $M$. If we have $1=(s_{1},\dots,s_{n})$ in $A$ and each $M_{s_{i}}$ is finite projective over $A_{s_{i}}$ then so is $M$.
Theorem 3.3 (Skolem-Noether).
If $A$ is an Azumaya algebra, of rank $\geqslant 1$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then $M=\{a\in A~|~\forall_{x}\,ax=\psi(x)a\}$ is a projective $k$-module of rank $1$.
Proof.
We use Theorem 3.1. The statement holds on the leaves ($M$ is even free of rank 11 on the leaves by Lemma 3.2) and descends to the root by tree induction using Theorem 3.2. ∎
If $k$ is local, the module $M$ is free of rank $1$ and any generator $u$ of $M$ is invertible (since it is invertible on the leaves) and $\psi$ is the inner automorphism defined by $u$, so we recover the usual form of the Skolem-Noether Theorem (any automorphism of a central simple algebra is an inner automorphism).
Claude Quitté has suggested the following refinement.
Theorem 3.4.
If $A$ is an Azumaya algebra, of rank $\geqslant 1$ as a projective module over a ring $k$, and $\psi$ is an automorphism of $A$, then
$M=\{a\in A~|~\forall_{x}\,ax=\psi(x)a\}$ and $N=\{a\in A~|~\forall_{x}\,a\psi(x)=xa\}$
are projective $k$-modules of rank $1$ and $M\otimes_{k}N\simeq k$ and $MN=k1$. Furthermore, $M$ and $N$ are direct factors in $A$.
Proof.
We have on the leaves $M_{S}N_{S}\subseteq S1$ and this statement descends to the root by tree induction. This can be used to define a map $M\rightarrow Hom(N,k),~x\mapsto(y\mapsto xy)$. This map is an isomorphism on the leaves, and this also descends to the root by tree induction. Finally, $M$ and $N$ are direct factors on the leaves, and this fact descends as well. ∎
4 Theory of algebraically closed local rings
To get the theory of algebraically closed local rings, we add to the theory of local rings the axioms
$\exists_{x}~P(x)=0$
for any monic nonconstant polynomial $P$.
Wraith conjectured in [Wraith] that this should define the classifying theory of the fppf topos. This question is discussed further in Blechschmidt’s PhD thesis [BT].
The same argument as in Lemma 3.1 shows the following result.
Lemma 4.1.
If $A$ is an Azumaya algebra over $R$, which is a local and residually discrete and algebraically closed $k$-algebra, then $A$ is a matrix algebra.
We can then prove as in the previous section
Theorem 4.1.
AA is an Azumaya algebra over $k$ if, and only if, we can build a finite tree of root $k$ using the following, where all leaves are matrix algebras:
  1. 1.
    $S$ is covered by $S\rightarrow S[1/u_{i}]$ if $1=(u_{1},\dots,u_{n})$;
  2. 2.
    $S$ is covered by $S\rightarrow S[X]/(P)$ if $P$ is monic nonconstant.
5 Conclusion
This paper is a beginning of the theory of Azumaya algebras in a constructive setting. The natural next step will be to analyse Gabber’s Theorem [KO] that the group of Azumaya algebras is the torsion part of the second cohomology group with coefficients in $GL_{n}(R)$ for the étale topology.